OpenStudy (anonymous):

1. What is the quotient

OpenStudy (anonymous):

see attachment

OpenStudy (anonymous):

help me plzzz..i have answer choices

OpenStudy (raden):

convert division 2 fractions be multiplication of them : a/b : c/d = a/b * d/c it is a start

OpenStudy (raden):

then factor out of them

OpenStudy (raden):

simplify, by cancel out for the like terms

OpenStudy (anonymous):

can yu give me an example...i dnt understand

OpenStudy (anonymous):

do yall want to see the answer choices

OpenStudy (anonymous):

start by factoring \[\frac{ (x+4)(x+4) }{ (2x-1)(x-4) }\div \frac{ (2x+6)(x+4) }{ 4x+4 }\]

OpenStudy (anonymous):

then change division to multiplication (multiply my reciprocal)\[\frac{ (x+4)(x-4) }{ (2x-1)(x-4) }*\frac{ 4x+4 }{ (2x+6)(x+4) }\]

OpenStudy (anonymous):

cancel out common terms \[\frac{ 4x+4 }{ (2x-1)(2x+6) }\]

OpenStudy (anonymous):

see pic for the animated and 3D process